OpenStudy (anonymous):

Find the quotient of quantity of 3 times x to the 3rd power plus 9 times x to the 2nd power minus 6 times x all over negative 3 times x. x^2 − 9x + 6 −x^2 − 3x − 2 x^2 + 9x − 6 −x^2 − 3x + 2

OpenStudy (anonymous):

\[(\frac{ 3x^3 + 9x^2 - 6x }{ -3x })\] \[= -x^2 - 3x + 2\]

OpenStudy (anonymous):

Just divide the term from each term above.

OpenStudy (anonymous):

Oh, wow, thanks so much :D